In a certain design, it is important that standard deviation of the current be less than 25 mA. In a test of n=18 parts, the sample standard deviation is found to be 20 mA. Can we be 95% confident that the standard deviation will not exceed 24 mA?
Answer #1

we have given,

sample standard deviation= 20 mA.

sample variance = 400mA

and n= 18

The critical values for α=0.05 and df = 17 degrees of freedom are χL2​=χ1−α/2,17​ =7.5642 and χU2​=χα/2,17 =30.191 (from chi square table )

95% confidence interval for variance = (n − 1) xổ 2,1m (n - 1)s xỉa/2-17

= 17*400 17 + 400、 7.5642:30.191 = (225.2326,898.9731)

95% confidence interval for standard deviation = (V225.2326,V898.9731​)=(15.0078,29.9829)

N0 ,We Cannot be 95% confident that the standard deviation will not exceed 24 mA